find the value of 7x+7 given that 2x-4=6
Answer by jim_thompson5910(35256)

2x-4=6


2x = 6+4


2x = 10


x = 10/2


x = 5


-------------------------------------------------------


7x+7


7(5)+7 ... replace x with 5 (since x = 5)


35 + 7


42


So the answer is 42